# Bumblebee gravity: spherically-symmetric solutions away from the potential minimum

### v1

@@ III.2.2 Spherical symmetry case, nonlinear solutions
Assuming spherical symmetry $B_{0}=B_{0}(r)$ we have for (18),
@@
This equation is a second order nonlinear ODE, which is nonautonomous (has explicit $x$ dependence). For possible potential choices we refer to Table 1. We begin studying the solution to this equation by noting several variable subsitutions that can be useful for studying different aspects of the solutions. Calling $2fV^{\prime}(-f^{2})=h(f)$, we consider two alternate forms:
@@
where the domain $0\leq x\leq\infty$ maps $1-1$ to the corresponding range for $u$ and $t$; namely $0\leq u\leq\infty$, $-\infty\leq t\leq\infty$, respectively.
To get an idea of the challenge of finding analytical solution consider the quadratic potential case, using (25):
$\ddot{f}+\alpha\ddot{f}=2\alpha^{2}\lambda e^{2\alpha t}\lambda(1-f^{2})f.$ (26)
@@ III.2.3 Stability Argument and Numerical Solutions
Using the substitution found in (24) can lead to a generalized statement about the behavior of the function for a given potential.
$u^{4}\frac{d^{2}f}{du^{2}}=2fV^{\prime}(-f^{2})$ (29)
@@
[…]
Equation (30) shows the stability argument for the quadratic potential, Eq. (31) applies the same reasoning to the Hypergeometric potential $V=g\Lambda^{4}[M(3,2,X/\Lambda^{2})-1]$, and Eq. (32) applies to $V=g\Lambda^{4}[M(4,2,X/\Lambda^{2})-1]$. Here, $f_{uu}$ denotes the second derivative of $f$ with respect to $u$, $\frac{d^{2}f}{du^{2}}$.
In the quadratic potential case, whenever $f\neq 1$ the concavity of $f(u)$ changes to move $f$ towards $f=1$ or $f=-1$. Thus $f_{uu}<0$ when $f>1$ and the $f_{uu}>0$ when $f<1$. This behavior occurs regardless of the initial position or initial slope $f_{u}$, and the solution asymptotes to one of the two points of stability $f=\pm 1$. The plot in Fig./ 2 shows how this argument applies regardless of the initial conditions.
Identical behavior occurs for the hypergeometric potential with $n=3$ case but the stable points are ±$\pm\sqrt{3}$. This is demonstrated in Fig. 3. However, with the Hypergeometric $n=4$ case, when $f<-\sqrt{6}$ or $f>\sqrt{6}$, the concavity of the solution moves it only further away from the u-axis, and thus it doesn’t stabilize, which is illustrated in Fig. 4. Overall then, we find that the first two potentials, quadratic and hypergeometric ($n=3$) are asymptotically stable, while $n=4$ is not.
Figure 2: Plot of a series of numerical solutions for the Quadratic potential $V=\frac{\lambda}{2}X^{2}$ with respect to $u=\frac{1}{x}$ with each solution showing dampening oscillations at ±$\pm 1$ as $u$ approaches 1, or $x$ approaches infinity
Figure 3: Plot of a series of numerical solutions for the Hypergeometric potential $n=3$, with each solution showing dampening oscillations around ±$\pm\sqrt{3}$
Figure 4: Plot of a series of numerical solutions for the Hypergeometric potential when $n=4$, with some solutions showing dampening oscillations around ±$\pm\sqrt{2}$, but solutions that move a distance of $\sqrt{6}$ from the origin escape completely as $u$ approaches 00, or as $x$ approaches infinity.
The argument used here applies when following decreasing values of $u$, which represents increasing values of $x$. What happens on the “other end”, when one approaches the origin. When looking at increasing values of $u$ or $x$ near the origin, we consider the following limit:
@@
where $h$ is the right-hand side of Eqns. (30-32). In the case of the Hypergeometric potentials, the function $g(f)$ is bounded, such that $|g(f)|\leq\epsilon$ for some positive constant $\epsilon$. We can then conclude that $f_{uu}$ approaches 00. This means that $f_{u}$ approaches some nonzero constant, such that the solution diverges as $u$ increases, or as $x$ approaches 00.
@@ IV Coupled gravity-vector solutions, Part I
This section contains a broad study of the field equations when gravity is present, confining attention to the effects of the potential $V$ as in (4). First we display the field equations for gravity and the bumblebee field. Variation of the action with (1), with respect to the metric $g_{\mu\nu}$ and the field $B_{\mu}$, yields
$G_{\mu\nu}$ $=\kappa\left[B^{\alpha}_{\phantom{\alpha}\mu}B_{\alpha\nu}-\frac{1}{4}g_{\mu\nu}B^{\alpha\beta}B_{\alpha\beta}-Vg_{\mu\nu}+2V^{\prime}B_{\mu}B_{\nu}\right],$ (34)
@@
where $G_{\mu\nu}$ is the Einstein tensor. These equation can be confirmed from Ref. [24] and others. Note that the antisymmetry of the field strength tensor $B_{\mu\nu}$ implies the current conservation identity
@@
We can make some general remarks on the weak and null energy conditions at this stage [101]. The first part of the stress-energy tensor (the first two terms on the right-hand side of the $G_{\mu\nu}$ part of (34)) is the standard electromagnetic piece, present in Maxwell theory. This can be shown to satisfy both the weak and null energy conditions, a widely known result. Thus, consider now the potential part $T_{V}^{\mu\nu}=-Vg^{\mu\nu}+2V^{\prime}B^{\mu}B^{\nu}$. Contraction on both indices with a unit timelike vector $u_{\mu}$ yields
@@
It is now unclear if the right-hand side is positive or zero, since the dependence is on $V^{\prime}$. Since $V$ is chosen to have a minimum for some value of $X$, $V^{\prime}$ can have positive and negative values away from that minimum. Thus, it appears there could be a violation of the null energy condition. Energy conditions in other vector models of spacetime-symmetry breaking have studied in [102].
@@
IV.1 Exact Anti de Sitter cosmological solutions
@@
where $\bar{V}$ is the potential evaluated at its minimum. This result then implies, via $G_{\mu\nu}=-\Lambda_{c}g_{\mu\nu}+...$, an effective cosmological constant
@@
Glancing at the potentials in Table 1, the only choices with nonzero minima for $V$ are the hypergeometric potentials. For example, for $V=g\Lambda^{4}[M(3,2,X/\Lambda^{2})-1]$, the two minima for $B_{0}$ occur when
@@
thus indicating a naturally occurring negative cosmological constant for the hypergeometric potentials. By natural we mean that the minimum value arises from the inherent nature of the potential itself, not by adding an ad-hoc constant to the action, like the usual $\Lambda_{c}$. This implies that Anti-de Sitter spacetime [104] can arise from the hypergeometric potentials in the bumblebee model. Indeed, the Friedman equations arising from (39) yield
@@
This type of solution (43) is an oscillating universe. Let us assume that the time scale is of the order of the age of the universe $\tau\approx 10Gyr\approx 3\times 10^{16}s$. Using the form of the (hypergeometric) potential above $\tau=5\times 10^{-43}s/(k\sqrt{g})$, where $k=\kappa\Lambda^{2}\sim\Lambda^{2}/M_{Pl}^{2}$ is the ratio of the mass scale of the bumblebee model to the Planck mass squared. To match the observed time scale of the universe we see that
@@ IV.2 Full equations and general features
$G_{tt}$ $=-\frac{AB^{\prime}}{B^{2}r}-\frac{A}{r^{2}}\left(1-\frac{1}{B}\right)=\kappa\left(\frac{B_{0}^{\prime 2}}{2B}-VA+2V^{\prime}B_{0}^{2}\right),$ (46)
@@
$G_{\theta\theta}$ $=\frac{r}{2B}\left(\frac{A^{\prime}}{A}-\frac{B^{\prime}}{B}\right)-\frac{r^{2}A^{\prime}}{4AB}\left(\frac{A^{\prime}}{A}+\frac{B^{\prime}}{B}\right)+\frac{r^{2}A^{\prime\prime}}{2BA}=-\kappa\left(\frac{r^{2}B_{0}^{\prime 2}}{2AB}+r^{2}V\right),$
@@
$B^{\prime}$ $=\frac{B}{r}(1-B)+\kappa\left(VB^{2}-\frac{BB_{0}^{\prime 2}}{2A}-2V^{\prime}\frac{B_{0}^{2}}{A}B^{2}\right)r,$
@@
The Reisner-Nordstrom solution is obtained for the case when $V=V^{\prime}=0$:
@@
where the $c_{n}$’s are fixed by the values of $B_{0}$, $B_{0}^{\prime}$, and $A$ at some position $r_{0}$. We will show below that exact solutions related to (50) are possible with a nonzero $V$.
@@ IV.3 Bekenstein horizon condition
It is known that for Brans-Dicke theory and the massive vector field case, as well as others, there are no static solutions with horizons except for the trivial case when the scalar or vector field is constant [107, 108, 109, 110]. Indeed, this is part of the “no-hair” conjecture, which says that stationary black holes are characterized only by their mass, charge, and angular momentum [111]. Exceptions to this theorem, while exotic and not corresponding to any standard model fields, abound [112, 113].
@@
$\int_{\cal V}d^{4}x\sqrt{g}\left(B_{\nu}\nabla_{\mu}B^{\mu\nu}-2V^{\prime}B_{\nu}B^{\nu}\right)=0.$ (51)
@@
$\int_{\partial\cal V}d^{3}\Sigma_{\nu}\sqrt{g}B_{\mu}B^{\mu\nu}-\int d^{4}x\sqrt{g}\left(\frac{1}{2}B_{\mu\nu}B^{\mu\nu}+2V^{\prime}B_{\nu}B^{\nu}\right)=0.$ (52)
At this stage one specializes the $4D$ volume ${\cal V}$ to be that bounded by two Cauchy (spacelike) hypersurfaces at different times, a portion of the black hole horizon, and a timelike 3-surface at spatial infinity.
The surface terms vanish due to the following: the integrals over the Cauchy hypersurfaces cancel each other when we assume a static spacetime; the asymptotic flatness assumption ensures that $B_{\mu\nu}\rightarrow 0$ at spatial infinity. The vanishing of the integral over the horizon is more subtle. The integral involves the scalar $d^{3}\Sigma_{\mu}\zeta^{\mu}$, where $\zeta^{\mu}=F^{\mu\nu}B_{\nu}$. Bekenstein has shown that this integral will vanish on the horizon provided that $\zeta^{\mu}$ is bounded on the horizon. We do not show the details here, the reader can refer to the original papers mentioned above.
It suffices for our work to argue that $\zeta_{\mu}\zeta^{\mu}$, which is spacelike on account of $F^{\mu\nu}B_{\mu}B_{\nu}=0$, is bounded on the horizon. Now, there should be no singularities of (physical) tensor quantities on the horizon, so $T_{\mu\nu}$ must be regular. For example, the trace $T_{\alpha}^{\phantom{\alpha}\alpha}=-4V+2V^{\prime}X$, where $X=g^{\mu\nu}B_{\mu}B_{\nu}=B_{0}^{2}/A$ with the spherical ansatz in (45). For the generic potential $V$, if $X$ is to be regular on the horizon then we must have $B_{0}=0$ on the horizon. Then then implies that $\zeta_{\mu}\zeta^{\mu}$ also vanishes on the horizon since it is proportional to $B_{0}^{2}$. Therefore $\zeta_{\mu}\zeta^{\mu}$ is bounded, and the horizon surface integral vanishes. This argument is similar to that used for the Einstein-Proca model in Ref. [108].
@@
The standard argument involves establishing in (53) that both terms in the integrand are positive definite and therefore must vanish since the integral over a spatial volume vanishes. For example, for the Proca vector field case with $V=\mu^{2}B_{\mu}B^{\mu}/2$, we get that $V^{\prime}=\mu^{2}/2\geq 0$ and so both terms in (53) must independently vanish. This then implies that $\partial_{r}B_{0}=0$ and thus the only way have a horizon in this model is the trivial case for $B_{0}$ that reduces to the Schwarzschild solution.

### v2

@@ III.2.2 Spherical symmetry case, nonlinear solutions
Assuming spherical symmetry ($B_{0}=B_{0}(r)$) we have for (18),
@@
This equation is an ordinary differential equation (ODE) that is second order, nonlinear, and nonautonomous (has explicit $x$ dependence). For possible potential choices we refer to Table 1. We begin studying the solution to this equation by noting several variable substitutions that can be useful for studying different aspects of the solutions. Calling $2fV^{\prime}(-f^{2})=h(f)$, we consider two alternate forms:
@@
where the domain $0\leq x\leq\infty$ maps 111-1 to the corresponding range for $u$ and $t$; namely $0\leq u\leq\infty$, $\infty\leq t\leq\infty$, respectively.
To get an idea of the challenge of finding analytical solutions consider the quadratic potential case, using (25):
$\ddot{f}+\alpha\dot{f}=2\alpha^{2}\lambda e^{2\alpha t}\lambda(1-f^{2})f.$ (26)
@@ III.2.3 Stability argument and numerical solutions
Using the substitution found in (24) can lead to a generalized statement about the behavior of the function for a given potential. The differential equation becomes,
$u^{4}\frac{d^{2}f}{du^{2}}=2fV^{\prime}(-f^{2}).$ (29)
@@
$u^{4}\frac{d^{2}f}{du^{2}}=2\lambda f(1-f^{2})\rightarrow\left\{\begin{array}[]{ll}f_{uu}>0&:\;f<-1\\ f_{uu}<0&:\;-1<f<0\\ f_{uu}>0&:\;0<f<1\\ f_{uu}<0&:\;f>1\\ \end{array}\right\},$ (30)
$u^{4}\frac{d^{2}f}{du^{2}}=ge^{-f^{2}}(3-f^{2})f\rightarrow\left\{\begin{array}[]{ll}f_{uu}>0&:f<-\sqrt{3}\\ f_{uu}<0&:-\sqrt{3}<f<0\\ f_{uu}>0&:0<f<\sqrt{3}\\ f_{uu}<0&:f>\sqrt{3}\\ \end{array}\right\},$ (31)
$u^{4}\frac{d^{2}f}{du^{2}}$ $=\frac{1}{6}ge^{-f^{2}}f\left(f^{4}-8f^{2}+12\right)$ (32)
$\rightarrow\left\{\begin{array}[]{ll}f_{uu}<0&:\;f<-\sqrt{6}\\ f_{uu}>0&:\;-\sqrt{6}<f<-\sqrt{2}\\ f_{uu}<0&:\;-\sqrt{2}<f<0\\ f_{uu}>0&:\;0<f<\sqrt{2}\\ f_{uu}<0&:\;\sqrt{2}<f<\sqrt{6}\\ f_{uu}>0&:\;f>\sqrt{6}\\ \end{array}\right\}.$
Equation (30) shows the stability argument for the quadratic potential, Eq. (31) applies the same reasoning to the hypergeometric potential $V=g\Lambda^{4}[M(3,2,z)-1]$, and Eq. (32) applies to $V=g\Lambda^{4}[M(4,2,z)-1]$, where $z=X/\Lambda^{2}$. Here, $f_{uu}$ denotes the second derivative of $f$ with respect to $u$, $\frac{d^{2}f}{du^{2}}$. Note that we can apply an argument about the trend of $f$ as $u$ tends toward smaller values because of the symmetry of the equations in (29) under the exchange $u\rightarrow-u$.
In the quadratic potential case, whenever $f\neq 1$ the concavity of $f(u)$ changes to move $f$ towards $f=1$ or f=1f=-1. Thus $f_{uu}<0$ when $f>1$ and the $f_{uu}>0$ when $f<1$. This behavior occurs regardless of the initial position or initial slope $f_{u}$, and the solution asymptotes to one of the two points of stability $f=\pm 1$. The plot in Fig. 2 shows how this argument applies regardless of the initial conditions.
Identical behavior occurs for the hypergeometric potential with the $n=3$ case but the stable points are $\pm\sqrt{3}$. This is demonstrated in Fig. 3. However, with the hypergeometric $n=4$ case, when $f<-\sqrt{6}$ or $f>\sqrt{6}$, the concavity of the solution moves it only further away from the u-axis, and thus it does not stabilize, which is illustrated in Fig. 4. Overall then, we find that the first two potentials, quadratic and hypergeometric ($n=3$) are asymptotically stable, while $n=4$ is not.
Figure 2: Plot of a series of numerical solutions for the quadratic potential $V=\frac{\lambda}{2}X^{2}$ with respect to $u=\frac{1}{x}$ with each solution showing dampening oscillations at $\pm 1$ as $u$ approaches 1, or $x$ approaches infinity. We set $\lambda=1$ for this plot. The colors represent distinct initial values for $f$ and $df/du$ at the point $u_{0}=1$.
Figure 3: Plot of a series of numerical solutions for the hypergeometric potential $n=3$, with each solution showing dampening oscillations around $\pm\sqrt{3}$. We set $g=1$ for this plot. The colors represent distinct initial values for $f$ and $df/du$ at the point $u_{0}=1$.
Figure 4: Plot of a series of numerical solutions for the hypergeometric potential when $n=4$, with some solutions showing dampening oscillations around $\pm\sqrt{2}$, but solutions that move a distance of $\sqrt{6}$ from the origin escape completely as $u$ approaches 0, or as $x$ approaches infinity. We set $g=1$ for this plot. The colors represent distinct initial values for $f$ and $df/du$ at the point $u_{0}=1$.
The argument used here applies when following decreasing values of $u$, which represents increasing values of $x$. What happens on the “other end”, when one approaches the origin? When looking at increasing values of $u$ or $x$ near the origin, we consider the following limit:
@@
where $h$ is the right-hand side of Eqns. (30-32). In the case of the hypergeometric potentials, the function $h(f)$ is bounded, due to the $e^{-f^{2}}$, such that $|h(f)|\leq\epsilon$ for some positive constant $\epsilon$. We can then conclude that $f_{uu}$ approaches 0. This means that $f_{u}$ approaches some constant, suggesting that the solution diverges as $u$ increases, or as $x$ approaches 0. This is similar to the electrostatic case with $V^{\prime}=0=V$ which has the exact solution $f=au+b$ for some constants $a$ and $b$.
@@ IV Coupled gravity-vector solutions, Part I
This section contains a broad study of the field equations when gravity is present, confining attention to the effects of the potential $V$ as in equation (4). First we display the field equations for gravity and the bumblebee field. Variation of the action in equation (1), with respect to the metric $g_{\mu\nu}$ and the field $B_{\mu}$, yields
$G_{\mu\nu}$ $=\kappa\Big{[}B^{\alpha}_{\phantom{\alpha}\mu}B_{\alpha\nu}-\frac{1}{4}g_{\mu\nu}B^{\alpha\beta}B_{\alpha\beta}$ (34)
$\phantom{space}-Vg_{\mu\nu}+2V^{\prime}B_{\mu}B_{\nu}\Big{]},$
@@
where $G_{\mu\nu}$ is the Einstein tensor. These equations can be confirmed from Ref. [24] and others. Note that the antisymmetry of the field strength tensor $B_{\mu\nu}$ implies the current conservation identity
@@
We can make some general remarks on the weak and null energy conditions at this stage [102]. The first part of the stress-energy tensor (the first two terms on the right-hand side of the $G_{\mu\nu}$ part of equations (34) is the standard electromagnetic piece, present in Maxwell theory. This can be shown to satisfy both the weak and null energy conditions, a widely known result. Thus, consider now the potential part $T_{V}^{\mu\nu}=-Vg^{\mu\nu}+2V^{\prime}B^{\mu}B^{\nu}$. Contraction on both indices with a unit timelike vector $u_{\mu}$ yields
@@
It is now unclear if the right-hand side is positive or zero, since the dependence is on $V^{\prime}$. Since $V$ is chosen to have a minimum for some value of $X$, $V^{\prime}$ can have positive and negative values away from that minimum. Thus, it appears there could be a violation of the null energy condition. Energy conditions in other vector models of spacetime-symmetry breaking have been studied in Ref. [103].
@@
IV.1 Exact anti-de Sitter cosmological solutions
@@
where $\bar{V}$ is the potential evaluated at its minimum. This result then implies, via $G_{\mu\nu}=-\Lambda_{c}g_{\mu\nu}+$..., an effective cosmological constant,
@@
Glancing at the potentials in Table 1, the only choices with nonzero minima for $V$ are the hypergeometric potentials. For example, for $V=g\Lambda^{4}[M(3,2,z)-1]$, the two minima for $B_{0}$ occur when
@@
thus indicating a naturally occurring negative cosmological constant for the hypergeometric potentials. By natural we mean that the minimum value arises from the inherent nature of the potential itself, not by adding an ad-hoc constant to the action, like the usual $\Lambda_{c}$. This implies that anti-de Sitter spacetime [105] can arise from the hypergeometric potentials in the bumblebee model. Indeed, the Friedmann equations arising from equation (39) yield
@@
This type of solution (43) is an oscillating universe. Let us assume that the time scale is of the order of the age of the universe $\tau\approx 10\,Gyr\approx 3\times 10^{17}s$. Using the form of the (hypergeometric) potential above, $\tau=5\times 10^{-43}s/(k\sqrt{g})$, where $k=\kappa\Lambda^{2}\sim\Lambda^{2}/M_{Pl}^{2}$ is the ratio of the mass scale of the bumblebee model to the Planck mass squared. To match the observed time scale of the universe we see that
@@ IV.2 Full equations and general features
$G_{tt}$ $=-\frac{AB^{\prime}}{B^{2}r}-\frac{A}{r^{2}}\left(1-\frac{1}{B}\right)$ (46)
$=\kappa\left(\frac{B_{0}^{\prime 2}}{2B}-VA+2V^{\prime}B_{0}^{2}\right),$
@@
$G_{\theta\theta}$ $=\frac{r}{2B}\left(\frac{A^{\prime}}{A}-\frac{B^{\prime}}{B}\right)-\frac{r^{2}A^{\prime}}{4AB}\left(\frac{A^{\prime}}{A}+\frac{B^{\prime}}{B}\right)+\frac{r^{2}A^{\prime\prime}}{2BA}$
$=-\kappa\left(\frac{r^{2}B_{0}^{\prime 2}}{2AB}+r^{2}V\right),$
@@
$B^{\prime}$ $=\frac{B}{r}(1-B)$
$\phantom{=}+\kappa\left(VB^{2}-\frac{BB_{0}^{\prime 2}}{2A}-2V^{\prime}\frac{B_{0}^{2}}{A}B^{2}\right)r,$
@@
The Reissner-Nordström (RN) solution is obtained for the case when $V=V^{\prime}=0$:
@@
where the $c_{n}$’s are fixed by the values of $B_{0}$, $B_{0}^{\prime}$, and $A$ at some position $r_{0}$. We will show below that exact solutions related to equation (50) are possible with a nonzero $V$.
@@ IV.3 Bekenstein horizon condition
It is known that for Brans-Dicke theory and the massive vector field case, as well as others, there are no static solutions with horizons except for the trivial case when the scalar or vector field is constant [108, 109, 110, 111]. Indeed, this is part of the “no-hair” conjecture, which says that stationary black holes are characterized only by their mass, charge, and angular momentum. Exceptions to this theorem, while exotic and not corresponding to any Standard Model fields, abound [112, 113].
@@
$\int_{\cal V}d^{4}x\sqrt{-g}\left(B_{\nu}\nabla_{\mu}B^{\mu\nu}-2V^{\prime}B_{\nu}B^{\nu}\right)=0.$ (51)
@@
$\int_{\partial\cal V}d^{3}\Sigma_{\nu}\sqrt{-g}B_{\mu}B^{\mu\nu}$ (52)
$-\int d^{4}x\sqrt{-g}\left(\frac{1}{2}B_{\mu\nu}B^{\mu\nu}+2V^{\prime}B_{\nu}B^{\nu}\right)=0.$
At this stage one specializes the four dimensional volume $\cal V$} to be that bounded by two Cauchy (spacelike) hypersurfaces at different times, a portion of the black hole horizon, and a timelike three-surface at spatial infinity.
The surface terms vanish due to the following: the integrals over the Cauchy hypersurfaces cancel each other when we assume a static spacetime; the asymptotic flatness assumption ensures that $B_{\mu\nu}\rightarrow 0$ at spatial infinity. The vanishing of the integral over the horizon is more subtle. The integral involves the scalar $d^{3}\Sigma_{\mu}\zeta^{\mu}$, where $\zeta^{\mu}=F^{\mu\nu}B_{\nu}$. Bekenstein has shown that this integral will vanish on the horizon provided that $\zeta^{\mu}$ is bounded on the horizon. We do not show the details here; the reader can refer to the original papers mentioned above.
It suffices for our work to argue that $\zeta_{\mu}\zeta^{\mu}$, which is spacelike on account of $F^{\mu\nu}B_{\mu}B_{\nu}=0$, is bounded on the horizon. Now, there should be no singularities of (physical) tensor quantities on the horizon, so $T_{\mu\nu}$ must be regular. For example, the trace $T_{\alpha}^{\phantom{\alpha}\alpha}=-4V+2V^{\prime}X$, where $X=g^{\mu\nu}B_{\mu}B_{\nu}=B_{0}^{2}/A$ with the spherical ansatz in (45). For the generic potential $V$, if $X$ is to be regular on the horizon then we must have $B_{0}=0$ on the horizon. This then implies that $\zeta_{\mu}\zeta^{\mu}$ also vanishes on the horizon since it is proportional to $B_{0}^{2}$. Therefore $\zeta_{\mu}\zeta^{\mu}$ is bounded, and the horizon surface integral vanishes. This argument is similar to that used for the Einstein-Proca model in Ref. [109].
@@
The standard argument involves establishing in equation (53) that both terms in the integrand are positive definite and therefore must vanish since the integral over a spatial volume vanishes. For example, for the Proca vector field case with $V=\mu^{2}B_{\mu}B^{\mu}/2$, we get that $V^{\prime}=\mu^{2}/2\geq 0$ and so both terms in (53) must independently vanish. This then implies that $\partial_{r}B_{0}=0$ and thus the only way have a horizon in this model is the trivial case for $B_{0}$ that reduces to the Schwarzschild solution.